1. Production possibility frontiers usually exhibit _________________ marginal ________________.

2. Anne can catch 10 fish in a day or pick 5 coconuts while Nancy can catch 12 fish or pick 9 coconuts. For each fish and coconuts, say who has an absolute advantage and who has a comparative advantage.

3. The gains from specialization are limited by the extent of ______________________.

4. We assume that production functions exhibit ___________________ marginal product of all inputs.

5. List the 3 categories of things that affect economic growth in the long run.

6. Which of the above categories probably explains most of the difference between rich countries and poor countries?

7. In the long run, money is ______________________________.

8. A _________________ conveys ownership of a company while a _________________ entitles the owner to a fixed amount of money in the future.

9. A riskier bond will have a ____________________ price and a ____________________ rate of return.

10. If output grows at 10% per year, how long will it take to double?

11. Countries experiencing “catching up growth” generally have __________________ growth rates and _______________________ output levels (per person) than those experiencing “cutting edge growth.”

II. Long Problems

1. An economy has the production function Y=AK1/2L1/2 where Y represents real output, K is capital, L is labor and A represents the level of technology. Let A=8, L=36, the savings rate be 1/4 and the depreciation rate be 1/5. Find the steady-state level of capital, output and consumption.

2. People become more patient (more willing to give up present consumption for future consumption)

a. Show the effect of this in the market for loanable funds and the bond market. (This requires two graphs.)

b. Show the effect on the steady-state levels of capital and output.

3. The following is a schedule giving the total product (Y) of a firm for different levels of investment (I).

I23456789101112

Y56.88.49.911.312.613.814.9516.0517.118.1

a. Find this firm’s demand schedule for investment depending on the real interest rate.

Let the supply of loanable goods be the following.

r0.05.10.15.2.25.3.35.4

S67891011121314

b. What is the equilibrium real interest rate and quantity of investment (use the above firm’s demand as the market demand). If the inflation rate is 0, what will the nominal rate be? What will the price of a one-year, risk-free bond with a face value of $200 be?

c. If the expected inflation rate rises to 3%, what will the nominal interest rate be and what will the price of the bond be? Show the effect of this increase in inflation on a graph of the loanable funds market and a graph of the bond market. (Two graphs.)
Answer #1

1. Production possibility frontiers usually exhibits diminishing marginal product.

2.

Fish Coconuts
Anne 10 5
Nancy 12 9

Nancy has absolute advantage in producing both goods because with given resources , Nancy can produce more fish and more coconuts than Anne.

Anne's Opportunity cost for producing fish = 5/10= 0.5 coconuts.

Nancy's opportunity cost for producing coconuts = 9/12 = 0.75 coconuts.

Because Anne has lower opportunity cost for producing fish , So, Anne has a comparative advantage in the production of fish. And Nancy has a comparative advantage in the production of coconuts.

3. The gains from specialization are limited by the extent of market.

4. We assume that production functions exhibit diminishing marginal product of all inputs.